OpenStudy (anonymous):

Solve each system of equations below. {y=x^2-2x-3 { {6x+y=2

OpenStudy (anonymous):

I substituted the one equation in for y and tried to solve, but my answer is in coordinates and I cant remember how to do that.

OpenStudy (campbell_st):

ok.. so let y = 2 - 6x substitute it into the 1st equation gives \[2 - 6x = x^2 - 2x - 3\] simplifying you get \[x^2 + 4x - 5 = 0\] this can be factorised to (x + 5)(x -1) = 0 should be easy to solve from here

OpenStudy (campbell_st):

you have a parabola and a straight line so there are up to 2 points of intersection... which seems to be the case in this question.

OpenStudy (anonymous):

so that gives me the x coordinates, how do you get the y?

OpenStudy (wikiemol):

plug the x back into either of the two equations.

OpenStudy (campbell_st):

substitute the values for x into the straight line 6x + y = 2

OpenStudy (anonymous):

Ah. that makes it pretty simple...thanks guys! I missed the factoring portion. :))

OpenStudy (campbell_st):

you can choose either equation but the straight line is the easier of the 2.